If weight of calcium containing 1mole of atoms is 40grams.Then what is the weight of carbon containing same number of atoms as in 10 grams of calcium.
1
Expert's answer
2019-06-26T04:03:32-0400

Solution.

If one mole of calcium contains 40 grams of calcium, then 10 grams:


"x = \frac{10 \times 1}{40}"

x = 0.25 mole


"m(C) = n(C) \times M(C)"

M(C) = 12 g/mole

m(C) = 0.25 * 12 = 3 g

Answer:

m(C) = 3 g